OpenStudy (anonymous):

3x2 – 6x = 17 solve this by completing the square

razor99 (razor99):

6x=3x2-17

razor99 (razor99):

6x=6-17 6x=-11 x=-11/6 x=1.83

razor99 (razor99):

thats the answer

razor99 (razor99):

helloo u there

OpenStudy (anonymous):

R u sure that's how you solve it by completing the square?